Q1. (25 marks)

Danish Inc purchased a computer on January 1, 2019, for $1,600,000. The straight-line method of depreciation was used, based on an expected life of 6 years and a residual value of $130,000. Prepare the journal entries to record depreciation for the first 6 months of 2021 and the sale of the computer on July 1, 2021, for $1,000,000. 
Answer #1

Depreciation per year = (cost-salvage value)/estimated life.

=(1600000-130000)/6

=$245,000 per year.

so Depreciation in 2019 is $245,000 and Depreciation in 2020 is $245,000

Depreciation first 6 months in 2021 is $245,000*6/12=$122,500

Total Depreciation as on july 2021 is $612,500.

Net book value on date of sale is $1,600,000-612,500 =$987,500.

journal entry in 2021

Accumulated Depreciaton A/C92 years DEP)........DR $490,000

Depreciation A/C (6 months Depreciation).. .....DR $122,500

TO computer A/c CR $612,500.

computer was sold for $1,000,000.

profit on sale :$1000,000-987500 = $12,500.

journa entry on date of sale .

Cash or bank A/c ........DR $1000,000.

To computer A/c .......CR $987,500

TO profit on sale ......CR $12,500.

being sale of computer for profit.